The 2018 balance sheet for Hallbrook Industries, Inc., is shown below. HALLBROOK INDUSTRIES, INC Balance Sheet December 31, 2818 (s in 8885) Assets Cash Short-term investments Accounts receivable Inventories Property, plant, and equipment (net) $348 298 348 33e 2,480 Total assets $3,78e Liabilities and Shareholders Equity Current liabilities Long-term liabilities Paid-in capital Retained earnings $ 540 498 1,450 1,228 Total liabilities and shareholders equity $3,78e The companys 2018 income statement reported the following amounts ($ in 000s): Net sales Interest expense Income tax expense $6,830 50 130 300 Net income Requlrec 1. Calculate the current ratio. (Round your answer to 2 declmal places.) 2. Calculate the acid-test ratio. (Round your answer to 3 decimal places.) 3. Calculate the debt to equity ratio. (R und your answer to 2 decimal places.) 4. Calculate the times interest earned ratio. (Round your answer to 1 declmel place.) 1 Current ratio 2. Acid-test ratio 3. Debt to equity ratio 4. Times interest earned ratio imes
Answer #1

Solution:

1.) Current ratio = Current Assets/Current Liabilities
=(340+290+340+330)/540
=2.41

2.) Acid-test ratio = (Cash + Short-term investments + Current Receivables)/Current Liabilities
=(340+290+340)/540
=1.80

3.) Debt to equity ratio = Total Liabilities/Total Equity
=(540+490)/(1440+1220)
=0.39

4.) Times interest earned ratio = Income before interest expense and taxes/Interest expense
=(300+50+130)/50
=9.6